### v3

# LLMs Meet Library Evolution: Evaluating Deprecated API Usage in LLM-based Code Completion

Abstract
Large language models (LLMs), pre-trained or fine-tuned on large code corpora, have shown effectiveness in generating code completions. However, in LLM-based code completion, LLMs may struggle to use correct and up-to-date Application Programming Interfaces (APIs) due to the rapid and continuous evolution of libraries. While existing studies have highlighted issues with predicting incorrect APIs, the specific problem of deprecated API usage in LLM-based code completion has not been thoroughly investigated.
To address this gap, we conducted the first evaluation study on deprecated API usage in LLM-based code completion. This study involved seven advanced LLMs, 145 API mappings from eight popular Python libraries, and 28,125 completion prompts. The study results reveal the status quo (i.e., API usage plausibility and deprecated usage rate) of deprecated API and replacing API usage in LLM-based code completion from the perspectives of model, prompt, and library, and indicate the root causes behind. Based on these findings, we propose two lightweight fixing approaches, ReplaceAPI and InsertPrompt, which can serve as baseline approaches for future research on mitigating deprecated API usage in LLM-based completion. Additionally, we provide implications for future research on integrating library evolution with LLM-driven software development.
I Introduction
Large language models (LLMs) [1, 2, 3, 4, 5, 6] have significantly advanced various aspects of software engineering [7], including code completion [8, 9, 10], code understanding [11, 12], defect detection [13, 14], and program repair [15, 16, 17]. These models, pre-trained or fine-tuned with extensive knowledge of code on large corpora, are effective for tailoring to different downstream tasks. In the realm of code completion, the state-of-the-art has evolved from statistics-based methods [18, 19] to LLM-based techniques [20, 21, 22]. Code completion is a sophisticated task that suggests variables, functions, classes, methods, and even entire code blocks, which depends on developers’ practical needs.
Motivation. To accelerate development, developers heavily rely on third-party libraries, interacting with them through Application Programming Interfaces (APIs). However, this reliance presents a challenge for code completion tools. Third-party libraries constantly evolve to undergo refactorings [23], fix bugs [24], apply security patches [25], or introduce new features. This rapid evolution leads to frequent API changes, with older APIs being deprecated and replaced by newer ones. Deprecated APIs are discouraged to use because of their incompatiblity with newer features or data, which will eventually disappear in future library updates [26]. Taking PyTorch [27], a popular deep learning library for instance, the API torch.gels() was deprecated in version 1.2 (August 2019) in favor of torch.lstsq(). Then, torch.lstsq() was deprecated in version 1.9 (June 2021) in favor of torch.linalg.lstsq(). Consequently, newly developed code should avoid using the deprecated torch.gels() and torch.lstsq(). Therefore, it’s crucial for code completion tools to suggest correct and up-to-dated APIs to developers.
Literature. However, to the best of our knowledge, the capabilities of LLM-based code completion regarding API deprecation is understudied [28, 29]. Although there emerges a substantial number of evaluation on code completion, a body of research focused on assessing the overall accuracy across various benchmarks [11, 30, 31, 32, 33]. Interestingly, Ding et al. [34] identified undefined names and unused variables as the most common syntactic errors produced by LLMs in Python code completions. Izadi et al. [35] found that incorrect function name predictions were prevalent, accounting for 23% of all token-level errors. Furthermore, Recent studies [36, 37] highlighted the issue of hallucinations in LLM-generated code. Their findings indicate the prevalence and potential risks of using unexpected APIs. Nevertheless, while researchers have noted the prevalence of incorrect function name predictions, they have not investigated this issue in depth. Library APIs, which constitute an important part in predicting external function names, are worth attached importance to.
Study. To address this gap, we conducted a study to examine the deprecated API usage in LLM-based code completion. The study aims to answer the primary research question:
What are the status quo (i.e., API Usage Plausibility (AUP) and Deprecated Usage Rate (DUR)) and potential root causes of deprecated and replacing API usage in LLM-based code completion?
This question is explored through three detailed aspects:
Model Perspective (RQ1) investigates the status quo and potential causes based on the performance of various LLMs;
Prompt Perspective (RQ2) examines the impact of different prompts on the status quo and potential causes;
Library Perspective (RQ3) analyzes the status quo and potential causes across different libraries.
To address these research questions, we conducted a series of experiments involving various libraries and LLMs. We collected 145 API mappings between deprecated APIs and their replacements from eight popular Python libraries. Based on these mappings, we retrieved 9,022 outdated functions and 19,103 up-to-dated functions using the deprecated APIs and replacing APIs, respectively. Then, we identify the locating lines of the deprecated or replacing APIs, mask them and subsequent code, and use the remaining parts as the code completion prompts. The original deprecated or replacing API is referred as the reference API. These prompts were then inputted into seven advanced code LLMs, including CodeLlama [5] and GPT-3.5, to generate completions and analyze the predicted API usages. If the predicted API usage corresponds to either the deprecated or replacement version of the reference API, it is annotated as plausible; otherwise, it is annotated as Others.
The study results reveal the following findings: Finding in RQ1: All evaluated LLMs encounter challenges in predicting plausible API usages and face issues with deprecated API usages, due to the presence of deprecated API usages during model training and the absence of API deprecation knowledge during model inference. Finding in RQ2: For the two categories of prompts derived from outdated and up-to-dated functions, the LLMs’ performance in predicting plausible and deprecated API usages differs significantly, influenced by the distinct code context characteristics of these prompts. Finding to RQ3: Across the eight libraries, the LLMs exhibit significant differences in their use of deprecated APIs, influenced by the characteristics of API deprecations during library evolution.
Lightweight Mitigation. Based on the study results and findings, we explored the feasibility of using two lightweight fixing approaches to mitigate deprecated API usage in LLM-based code completion. Given a completion containing a deprecated API usage, the first approach, named ReplaceAPI, directly replaces the deprecated API usage with the replacement and regenerates the remaining parts (e.g., argument list) during the decoding process. The second approach, named InsertPrompt, inserts an additional replacing prompt after the original prompt to guide the LLMs to use the replacing API and then regenerate the completions. We then evaluate the effectiveness of the proposed approaches in terms of fixing deprecated API usages and the accuracy in predicting line-level completions (RQ4). The evaluation results demonstrate that ReplaceAPI effectively addresses deprecated API usages for all evaluated open-source LLMs, achieving fix rates exceeding 85% with acceptable accuracy measured by Edit Similarity and Exact Match compared to ground-truth completions. While InsertPrompt does not currently achieve sufficient effectiveness and accuracy in fixing completions containing deprecated API usage, it shows potential for future exploration.
Contribution. This paper makes the following contributions:
  • We conducted the first study that reveals the status quo and causes of deprecated API usages in LLM-based code completion from prespectives of model, prompt, and library.
  • We proposed two lightweight approaches, named ReplaceAPI and InsertPrompt, to serve as baselines for mitigating deprecated API usage in LLM-based completion.
  • We provide implications for future research on the synergy of library evolution and LLM-driven software development.
II Related Work
II-A Library Evolution
Library evolution involves refactorings [23], bug fixes [24], and new feature introductions. Typically, refactorings can deprecate old APIs and introduce new replacements. Several studies have examined the reasons that developers deprecate APIs and how the clients react to such deprecations [38, 39, 40, 41]. The reasons include improving readability, reducing redundancy, avoiding bad code practices and fixing functional bugs. Deprecated APIs can affect hundreds of clients [42], particularly when clients struggle to keep pace with rapidly evolving software [43, 44]. McDonnell et al. [45] found that only 22% of outdated API usages are eventually upgraded to use replacing APIs. Similarly, Hora et al. [46] found that client developers consumes considerate time to discover and apply replacing APIs, with the majority of systems not reacting at all. When clients do not upgrade their APIs, they silently accumulate technical debt in the form of future API changes when they finally upgrade [47]. To locate the replacing API, existing works leverage change rules written by developers [48], developer recordings [49], similarity matching [50], mining API usage in libraries [51], and in client projects [52]. Henkel and Diwan [49] developed an IDE plugin that allows library developers to record API refactoring actions and client developers to replay them. Godfrey and Zou [53] proposed a semi-automated origin analysis using similarities in name, declaration, complexity metrics, and call dependencies. Wu et al. [54] introduced a hybrid approach combining call dependency and text similarity analysis to identify API change rules. Recently, [55] proposed RepFinder to find replacing APIs for deprecated APIs in library updates from multiple sources.
In this work, we aim to comprehend the statuses and causes of deprecated API usage in LLM-based code completion and provide implications for mitigating the deprecated API usages.
II-B LLM-based Code Completion
Code completion is an important functionality in modern IDEs and editors. Historically, researchers have explored statistical models [18, 19]. With the advent in natural language processing, researchers have embraced deep learning for code completion [8, 9] because they are similar in token-based prediction. To explore the capability of code completion tools driven by LLMs [20, 21, 22], numerous evaluations of LLMs have been proposed. Ciniselli et al. [32, 33] conducted a large-scale study exploring the accuracies of state-of-the-art Transformer-based models in supporting code completion at various granularity levels, from single tokens to entire code segments. Zeng et al. [11] found that pre-trained models significantly outperform non-pre-trained state-of-the-art techniques in program understanding tasks. They also reveal that no single pre-trained model dominates across all tasks. Xu et al. [30] evaluated the performance of LLMs on the HumanEval dataset. Ding et al. [34] identified undefined names and unused variables as the most common errors produced by language models in Python code completions. Izadi et al. [35] evaluated the LLMs using real auto-completion usage data across 12 languages. They found that incorrect function name predictions, were prevalent, accounting for 23% of all token-level errors. Besides, Liu et al. [31] proposed EvaluPlus, which benchmarks the functional correctness of LLM-synthesized code using test cases. In addition to accuracy concerns, LLM-based approaches face issues such as security vulnerabilities and hallucinations. Sallou et al. [56] explored threats posed by LLMs, including unpredictability in model evolution, data leakage, and reproducibility. Liu et al. [36] categorized the hallucinations brought by LLM-generated code.
The findings on incorrect function predictions partially motivate our study. However, we focus on the severity of predicting deprecated API usages in LLM-based code completion.
III Study Setup
Figure 1: Overview of Our Study
We chose Python, a popular programming language which ranks first among the most popular ones based in the recent year [57]. We targeted eight popular Python libraries. Five of these libraries were used in a previous study on Python API deprecation [25], including Numpy, Pandas, scikit-learn, SciPy, and seaborn. Additionally, we added three popular deep learning libraries, i.e., TensorFlow, PyTorch, and Transformers. The setup of our study is presented in Figure 1. It includes four steps. The API Mapping Collection gathers mappings between deprecated APIs and their replacements from various libraries. The Completion Prompt Construction step involves creating completion prompts by identifying instances of deprecated and replacing API usage in open-source Python repositories. The LLM-Based Code Completion step uses various LLMs to generate code completions for these prompts. Finally, the Completion Result Annotation step automatically annotates the generated completions and calculates relevant metrics.
III-A API Mapping Collection
We identified API mappings (i.e., deprecated APIs and the mapping replacements) from the documentation and change logs from each library version following the previous study [25]. For each library, we selected their versions that were released after January 2019. We chose January 2019 as the starting point for collecting API mappings, providing a five-year window that balances sufficient version numbers with manageable human effort. The version ranges are presented in Table I. Specifically, we reviewed the documentation and change logs of each library and manually look for deprecated API occurrences which indicate the corresponding the mapping replacements. For instance, in the API documentation of PyTorch, version 1.9.0 [58], a deprecation message indicates that “torch.lstsq() is deprecated in favor of torch.linalg.lstsq() and will be removed in a future PyTorch release.”, where the mapping of the deprecated API to the replacing API is torch.lstsq $\to$ torch.linalg.lstsq. For one-to-many mappings (i.e., one deprecated API mapped to multiple replacing APIs), we split them into many one-to-one mappings. Two authors independently collected the data, with a third author resolving inconsistencies. The process took about three working days, and the Jaccard coefficient between the first two collectors’ mappings was 91.8%. In total, we collected a preliminary number of 247 API mappings.
TABLE I: Statistics of our Collected API Mappings
  Library Version Range # Mappings # Functions
Outdated Up-to-dated
  Numpy 1.16.0-1.26.4 3 567 2,988
Pandas 0.24.0-2.2.2 10 69 69
scikit-learn 0.21.3-1.5.0 18 985 1,197
SciPy 1.2.1-1.13.0 4 245 1,458
seaborn 0.9.1-0.13.2 3 904 1,329
TensorFlow 1.13.1-2.16.1 57 1,491 4,830
PyTorch 1.0.1-2.3.0 21 4,726 6,406
Transformers 1.0.0-4.40.2 29 100 63
  Total 145 9,022 19,103
III-B Completion Prompt Construction
We constructed code completion prompts by searching real-world code snippets that contain usages of either deprecated APIs or replacing APIs from collected API mappings in open-source Python repositories, to simulate realistic code completion scenarios.
III-B1 Outdated and Up-to-dated Function Location.
We utilized Sourcegraph [59], a widely used code search service to search code snippets. It supports integration with GitHub where we can retrieve Python source files from millions of open-source code repositories. For each deprecated or replacing API, we constructed search queries using both its full qualified name (FQN) (e.g., torch.lstsq) and a logical disjunction of its constituent parts (e.g., “torch AND linalg AND lstsq”) to ensure comprehensive retrieval. For each retrieved Python source file, we parsed it into an Abstract Syntax Tree (AST) and extracted the containing functions that invoked the deprecated or replacing APIs. Specifically, we located function definition nodes in the AST and traversed its descendants. For each descendant, we checked if it is a function call node and matched the function call to the deprecated or replacing APIs. To correctly match the function call via API FQNs, we performed lightweight object type resolution and alias resolution, similar to [25].
  • Object Type Resolution: In the object-oriented programming (OOP) languages, the APIs can be encapsulated into a class as a method. Therefore, determining the FQN of the API invocation need to resolve the corresponding type of the invoking object. For example, the pandas library defines a core class DataFrame with a member method loc() and the client creates an object of class DataFrame, assigns to a variable dt, and invokes the method using dt.loc(). Typically, it requires resolving the type of dt. To that end, we analyzed the assign statements to track object definitions, enabling us to determine the class names for objects in function calls and infer the called APIs. For instance, if the object “dt” in the call dt.loc() was created in a preceding assign statement (dt = pandas.DataFrame(…)), we could infer that the corresponding API was pandas.DataFrame.loc().
  • Alias Resolution: Developers can alias packages, classes, and functions in Python using the import-as feature [60]. This mechanism requires resolving API aliases by analyzing import statements. For example, the pandas package is often imported with the alias “pd” via the statement import pandas as pd. In this case, pd.DataFrame.loc() was resolved to pandas.DataFrame.loc(). Additionally, Python provides a from-import mechanism allowing developers to use APIs with short names instead of their FQNs. For example, through from torch.linalg import lstsq, the API in torch.linalg can be directly called via lstsq(). These short names were resolved by analyzing the import statements.
After the lightweight object type resolution and alias resolution, we obtained the corresponding FQN for each function call. We checked whether the corresponding FQNs matched the APIs in the collected API mappings, identifying the first matched API as the reference API.
We denote the containing function as an outdated function if a deprecated API was matched. Meanwhile, we denote it as an up-to-dated function if a replacing API was matched. We collected 113,660 Python source files by querying SourceGraph. We filtered out API mappings that returned either no instances of outdated or up-to-dated functions. As a result, we collected 9,022 outdated and 19,103 up-to-dated functions from 145 API mappings. The statistics are presented in Table I.
Figure 2: Illustration of Completion Prompt Construction for An Up-to-dated Function.
III-B2 Incomplete Code Extraction.
In the task of code completion, developers usually have started with a few lines of code and pause in the middle, waiting for LLMs to return the suggested content based on the upward context. Therefore, to evaluate the performance of LLMs in the scenario, we constructed the line-level code completion prompts. For each outdated or up-to-dated function, we located the invocation line of the deprecated or replacing APIs, respectively. We collected the preceding lines before the invocation line into our line-level code completion prompts for an outdated or up-to-dated function, which is usually incomplete.
Figure 2 represents one of our collected up-to-dated functions. The function invokes a API of PyTorch, i.e., torch.linalg.lstsq in the fourth line. The line-level code completion prompt for this function is highlighted in the wine-red dotted rectangle.
After processing all outdated and up-to-dated functions, we obtained two corresponding datasets, denoted as $\mathcal{O}$ and $\mathcal{U}$, respectively. Each sample in $\mathcal{O}$ and $\mathcal{U}$ was formatted as ($pmpt$, $dep\to rep$), where $pmpt$ is a code completion prompt $pmpt$ and $dep\to rep$ denotes an API mapping from the deprecated API to the corresponding replacing API.
III-C LLM-based Code Completion
We leverage multiple LLMs including open-source and closed-soure with varying parameter sizes and observe their performance on the code completion task. The complete LLM list is presented in Table II.
  • CodeGen-350m, 2b, 6b: CodeGen [3] is a family of LLMs developed by Salesforce specifically for code generation.
  • DeepSeek-1.3b: DeepSeek-Coder [6] is designed on top of Transformer and tailored for code-related applications.
  • StarCoder2-3b: StarCoder2 [4] is an LLM optimized for coding tasks. It leverages large-scale pre-training on code datasets to understand programming languages deeply.
  • CodeLlama-7b: CodeLlama [5] is a specialized variant of Meta’s LLaMA [61], adapted for programming tasks.
  • GPT-3.5: GPT-3.5 [62] is a general-purpose language model developed by OpenAI. While it is not exclusively designed for coding, it possesses powerful code generation capabilities due to its extensive training on diverse text, including programming languages.
The first six code LLMs were downloaded from Hugging Face [63]. For LLMs with Python-specific versions available, we utilized those versions to improve completion results for Python functions. Consequently, the versions used for CodeGen and CodeLlama were CodeGen-{350M,2B,6B}-mono and CodeLlama-7b-Python-hf, which were fine-tuned on additional Python corpora. For DeepSeek and StarCoder2, the versions employed were deepseek-coder-1.3b-instructstarcoder2-3b, respectively. For the general purpose LLM, i.e., GPT-3.5, we queried the model via its official online APIs [64], using the gpt-3.5-turbo version released in January 2024.
For the six LLMs specifically tailored for code-related tasks, i.e., three versions of CodeGen, DeepSeek, StarCoder2, and CodeLlama, the constructed prompts can be directly fed into the models to generate completions. Meanwhile, for the general-purpose GPT-3.5, we provided an instruction to specify the task, preventing the model from performing beyond code completion. The instruction was: “Complete and output the next line for the following Python function: $pmpt$. For all these LLMs, we utilized greedy decoding (i.e., choosing the token with highest possibility at each decoding step) to generate one completion for each prompt in $\mathcal{O}$ or $\mathcal{U}$. The maximal output token limit was set to 50. The greedy decoding for GPT-3.5 was implemented through setting the temperature parameter to 0.
The procedure of LLM-based completion is defined as:
$comp\leftarrow\textbf{LLM}(pmpt)$
TABLE II: Evaluated Large Language Models (LLMs)
  Model # Params Python Fine-Tuned Open-Source
  CodeGen-350m 350 M
CodeGen-2b 2 B
CodeGen-6b 6 B
DeepSeek-1.3b 1.3 B
StarCoder2-3b 3 B
CodeLlama-7b 7 B
GPT-3.5 175 B UNKNOWN
III-D Completion Result Annotation
For each sample ($pmpt$, $dep\to rep$), we examined the completions generated by LLMs and determine whether the studied API was predicted and whether the deprecated API or replacing API was predicted. Specifically, we extract the FQN of the API invocation in the predicted line using the same object type resolution and alias resolution in Sec. III-B1. The annotation procedure is formally described as follows:
$\{\emph{DepC},\emph{RepC},\emph{Others}\}\leftarrow\text{anno}(comp)$
Specifically, DepC denotes that the LLM gives a Deprecated Completion suggestion, and RepC denotes that the LLM gives a Replacing Completion suggestion. We identify DepC and RepC by matching the FQN of an invocating API to either $dep$ or $rep$. Others denotes that the LLM suggests neither of the mapping APIs. Moreover, if a completion was annotated as either DepC or RepC, we treat it as plausible. This indicates that the LLM successfully understood the code context and selected a plausible API functionality.
III-E Metrics
We investigate the performance of the LLMs using the following metrics:
  • API Usage Plausibility (AUP): This metric measures the portion of plausible completions, which were annotated as DepC or RepC. AUP is defined as:
    $AUP=\frac{1}{|\mathcal{P}|}\sum_{p\in\mathcal{P}}\mathbb{I}(\text{anno}(\textbf{LLM}(p))\in\{\emph{DepC},\emph{RepC}\})$
  • Deprecated Usage Rate (DUR): This metric calculates the rate of plausible completions that were annotated as DepC. DUR is defined as:
    $DUR=\frac{\sum_{p\in\mathcal{P}}\mathbb{I}(\text{anno}(\textbf{LLM}(p))=\emph{DepC})}{\sum_{p\in\mathcal{P}}\mathbb{I}(\text{anno}(\textbf{LLM}(p))\in\{\emph{DepC},\emph{RepC}\})}$
$\mathcal{P}$ is the prompt set (i.e., $\mathcal{O}$ or $\mathcal{U}$), and $\mathbb{I}(\cdot)$ is a binary function that returns 1 if the passed argument is true and 0 otherwise.
The API Usage Plausibility (AUP) measures how effectively LLMs predict “accurate” APIs without considering their deprecation status. This metric is crucial because merely counting deprecated API usages (DUR) does not fully capture how LLMs manage API deprecations. For instance, a low number of deprecated API usages might be due to the prevalence of Others completions, suggesting influences beyond deprecation. To provide fair comparisons across models, prompts, and libraries, we calculate the Deprecated Usage Rate (DUR) based on AUP. A balanced relationship between AUP and DUR, characterized by a relatively high AUP and low DUR, indicates that LLMs effectively predict both up-to-date and “accurate” APIs. In contrast, an imbalance suggests either a lower rate of “accurate” predictions or a prevalence of outdated APIs. All the LLMs exhibited low AUPs due to the generation of many Others predictions, likely because many APIs (e.g., PyTorch APIs) share similar usage contexts and offer flexible combinations. In addition, although these Others may provide alternative solutions for similar functionalities, they fall outside our focus on LLMs’ handling of API deprecations. To clarify, we formally defined both Others and AUP to avoid confusion with terms like “wrong” or “incorrect”. Nevertheless, we recognize the value of exploring this issue and leave a broader question for future work: how to effectively evaluate the functional correctness of generated completions, given the numerous alternative implementations for the same goal.
IV Study Results
We present the experimental results and key findings on the status quo and root causes of deprecated and replacement API usage in LLM-based code completion. As illustrated in Figure 1, the results are categorized into three detailed aspects: Model Perspective (RQ1), Prompt Perspective (RQ2), and Library Perspective (RQ3). For each RQ, the results and findings are presented by first showing the status quo through API Usage Plausibility and Deprecated Usage Rate, followed by providing an in-depth Root Cause Analysis.
IV-A RQ1: Model Perspective for Status Quo and Root Causes
IV-A1 Status Quo Analysis
Figure 3 shows the distribution of RepC and DepC completions by the different LLMs, and Table III presents the AUP and DUR metrics.
API Usage Plausibility. The completion distribution and the low AUP highlight that all the LLMs faced challenges in predicting plausible API usages for the given prompts. The AUP of the LLMs for overall dataset (i.e., All $=\mathcal{O}\cup\mathcal{U}$) ranges from 9% to 23%, indicating that a majority of predictions were Others. Among the LLMs, CodeLlama-7b achieved the highest AUP (23.4%), while StarCoder2-3b had the lowest overall AUP (9.3%). This may be attributed to the fact that StarCoder2 was not specifically fine-tuned on additional Python corpora, unlike other LLMs such as CodeLlama and CodeGen. Comparing the three versions of CodeGen suggests that the capacity of LLMs to predict plausible API usages increased with model size (i.e., 11.0%, 28.8%, and 23.6% for CodeGen-350m, -2b, and -6b, respectively), given the model architecture and training data remain consistent. However, it’s noteworthy that the largest LLM, GPT-3.5, did not achieve a high AUP (13.2%), possibly due to the instruction used not being finely tuned with advanced prompt engineering techniques such as chain-of-thought (COT) [65] and in-context learning (ICL) [66].
Summary 1: All the evaluated LLMs faced challenges in predicting plausible API usages, with AUP ranging from 10% to 30%. Effectiveness of LLM-based code completion generally improved with model size and language-specific fine-tuning.
Figure 3: Distribution of RepC, DepC, and Others Completions by LLMs for Prompts from $\mathcal{O}$ and $\mathcal{U}$
TABLE III: AUP and DUR for $\mathcal{O}$, $\mathcal{U}$, and $\mathcal{O}\cup\mathcal{U}$
  Model AUP (%) DUR (%)
$\mathcal{O}$ $\mathcal{U}$ All $\mathcal{O}$ $\mathcal{U}$ All
  CodeGen-350m 7.9 12.4 11.0 77.7 9.0 24.9
CodeGen-2b 15.9 20.1 18.8 89.6 11.3 32.6
CodeGen-6b 20.7 23.6 22.7 90.0 11.1 34.2
DeepSeek-1.3b 9.2 11.8 11.0 69.7 11.6 27.2
StarCoder-3b 8.1 9.9 9.3 79.4 17.0 34.4
CodeLlama-7b 21.6 24.3 23.4 86.9 12.1 34.2
GPT-3.5 11.9 13.8 13.2 85.7 17.8 37.4
Deprecated Usage Rate. The distribution shown in Figure 3, along with the DUR metric presented in Table III, indicates that all LLMs faced issues with using deprecated API usages.
The DUR of the LLMs for the overall dataset (i.e., All $=\mathcal{O}\cup\mathcal{U}$) ranges from 25% to 38%, with larger models (e.g., CodeGen-6b, CodeLlama-7b, and GPT-3.5) generally predicting more usages of deprecated APIs. Considering the differences among the LLMs, CodeLlama-7b and CodeGen-6b demonstrated the best balance between AUP and DUR. They achieved significant improvements in AUP compared to other LLMs, with a comparable DUR of 34.2%. Conversely, StarCoder2-3b and GPT-3.5 exhibited higher DUR (i.e., 34.4% and 37.4%) despite having much lower AUP. These results indicate that the preference of LLMs for using deprecated or replacing APIs is not closely related to their capacity for predicting plausible completions.
Summary 2: All the evaluated LLMs faced issues with deprecated API usages, with DUR ranging from 25% to 38%, and larger models exhibiting higher DUR. Among the LLMs, CodeLlama-7b and CodeGen-6b demonstrated the best balance between AUP and DUR.
IV-A2 Potential Cause Discussion
From the model perspective, the causes of deprecated API usages consist of two main points:
Model Training: As libraries evolve, both deprecated APIs and their replacements are commonly found in open-source code repositories. Since the training data for LLMs are primarily sourced from these repositories without filtering for deprecated APIs, they often include instances of deprecated API usage. While the training datasets for specific LLMs like CodeLlama and GPT-3.5 are not publicly available, preventing direct inspection, the distribution of deprecated API usages in open-source code repositories can serve as a proxy for understanding the nature of their training data. In this study, we identified 9,022 instances of deprecated API usage and 19,103 instances of their replacements in open-source code repositories. These coarse statistics suggest that a significant portion of the training data likely includes deprecated APIs. Moreover, prior studies have demonstrated a relationship between training data distribution and model output behavior [67, 68]. When trained on data containing deprecated APIs, LLMs may “memorize” these APIs and associated usage patterns as part of their learned knowledge [69, 70]. The different training datasets also led to different AUP and DUR of the LLMs.
Model Inference: LLMs generated completions by predicting token probabilities based on their learned prior knowledge (e.g., memorized API usage contexts) and applying token selection strategies (e.g., greedy search or beam search). Given certain contexts, LLMs were likely to predict deprecated API usages due to high token probabilities, without considering any posterior API deprecation knowledge. To support this hypothesis, we calculated the generative likelihoods of both $dep$ and $rep$ for each prompt $pmpt$ across open-source LLMs by analyzing predicted token probabilities during decoding. To begin, we input $pmpt$ into an LLM, followed by sequentially feeding each token in $dep$ (or $rep$) into the model. Throughout this process, we recorded the predicted token probability distributions prior to each token in $dep$ (or $rep$) being input. Next, we extracted the predicted probabilities for tokens in $dep$ (or $rep$) from these distributions using their vocabulary indices. The generative likelihoods of $dep$ and $rep$ were then computed by summing the log probabilities of their tokens, a common method in language modeling [71]. We performed a paired t-test on these likelihoods. The t-statistic values indicate that for all six open-source LLMs, the likelihoods of deprecated APIs are significantly higher than those of replacing APIs for prompts from $\mathcal{O}$, whereas the opposite is true for prompts from $\mathcal{U}$. The $p$ values are approximately zero.
Summary 3: There are two primary reasons why LLMs predict deprecated APIs: the presence of deprecated API usages in corpora during model training, and the absence of posterior knowledge about API deprecations during model inference.
IV-B RQ2: Prompt Perspective for Status Quo and Root Causes
IV-B1 Status Quo Analysis
Table III presents the AUP and DUR for prompts from the two datasets, i.e., $\mathcal{O}$ and $\mathcal{U}$.
API Usage Plausibility. Between the prompts from the two different datasets, $\mathcal{O}$ and $\mathcal{U}$, there are some differences in AUP for all LLMs, with relative differences (i.e., $(AUP^{\mathcal{U}}-AUP^{\mathcal{O}})/AUP^{\mathcal{O}}$) ranging from 12.5% to 60.0%. This disparity may be attributed to the imbalance in the number of outdated and up-to-dated functions in the LLMs’ training corpora [72, 73]. Indirect evidence for this is that, in this study, the up-to-dated functions collected from open-source code repositories were about twice as many as the outdated functions (i.e., 19,103 vs. 9,022), even though there was no collection preference. Given that LLMs were often trained on open-source code repositories, they likely learned more up-to-dated functions than outdated functions, leading to better AUP for the $\mathcal{U}$ dataset. Additionally, larger LLMs showcased smaller AUP differences (e.g., 16% for CodeLlama-7b), possibly due to the better generalizability.
Summary 4: The LLMs showcased difference in AUP between the two datasets $\mathcal{O}$ and $\mathcal{U}$. This difference is possibly attributed to the different distribution of outdated and up-to-dated functions in the training corpora of LLMs.
Deprecated Usage Rate. When considering $\mathcal{O}$ and $\mathcal{U}$ separately, all LLMs consistently demonstrated extremely high deprecated usage rates for $\mathcal{O}$ (70%-90% DUR) and relatively low rates for $\mathcal{U}$ (9%-18% DUR). This significant difference is also evident in the distribution of RepC and DepC completions shown in Figure 3. This observed discrepancy can be attributed to the tendency of LLMs to predict reference APIs (i.e., the APIs used in the original functions) for most prompts in both the $\mathcal{O}$ and $\mathcal{U}$ datasets, reflecting the unique contextual characteristics of each prompt. To quantify this, we define an auxiliary metric, the Reference Usage Rate (RUR), as follows:
$\text{RUR}=\frac{\sum_{p\in\mathcal{P}}\mathbb{I}(\textbf{LLM}(p)=reference)}{\sum_{p\in\mathcal{P}}\mathbb{I}(\text{anno}(\textbf{LLM}(p))\in\{\text{\emph{DepC}},\text{\emph{RepC}}\})},$
which measures the proportion of reference APIs in plausible completions. RUR is equal to DUR for $\mathcal{O}$ and $(1-DUR)$ for $\mathcal{U}$. Thus, no significant difference exists between $\mathcal{O}$ and $\mathcal{U}$ in terms of RUR, which is consistently within the 70%-90% range for both. Nonetheless, the 9%-18% DUR for $\mathcal{U}$ indicates that LLMs still predicted the usage of deprecated APIs, even for the prompts from up-to-dated functions.
Summary 5: The LLMs consistently exhibited a significant difference in DUR between the two datasets, with extremely high deprecated API usage rates for $\mathcal{O}$ (70%-90% DUR) and relatively low rates for $\mathcal{U}$ (9%-18% DUR).
IV-B2 Potential Cause Discussion
The contextual characteristics of the input completion prompts can significantly influence LLMs’ use of deprecated APIs. Since the completions were generated based on the input prompts, specific contexts can lead LLMs to use deprecated APIs. As presented above, the LLMs showed significantly different DUR for prompts from $\mathcal{O}$ and $\mathcal{U}$. Upon comparing the prompts from the two datasets, we found that contextual characteristics of the prompts, such as specific variables and function calls, lead to the discrepancy of LLMs’ predictions, i.e., whether to use deprecated APIs or their replacements.
To quantitatively assess differences in contextual characteristics, we conducted lightweight contextual feature extraction for prompts from $\mathcal{O}$ and $\mathcal{U}$ using static analysis. Specifically, we grouped prompts according to their reference API, denoted as $api$, with each group represented as $\mathcal{P}(api)$. For each prompt in $\mathcal{P}(api)$, we identified the function used to construct it (cf. Section III-B) and extracted the variable and function call names in this function. After analyzing all prompts within $\mathcal{P}(api)$, the extracted variable and function call names were compiled into a contextual feature set, represented by $\mathbf{Feat}(\mathcal{P}(api))$, for the prompt group. Through this approach, for each API mapping $dep\to rep$, we obtained two paired prompt groups $\mathcal{P}(dep)\subset\mathcal{O}$ and $\mathcal{P}(rep)\subset\mathcal{U}$, along with their feature sets, $\mathbf{Feat}(\mathcal{P}(dep))$ and $\mathbf{Feat}(\mathcal{P}(rep))$. We then assessed the contextual similarity between $\mathcal{P}(dep)$ and $\mathcal{P}(rep)$ using $\frac{|\mathbf{Feat}(\mathcal{P}(dep))\cap\mathbf{Feat}(\mathcal{P}(rep))|}{|\mathbf{Feat}(\mathcal{P}(dep))\cup\mathbf{Feat}(\mathcal{P}(rep))|}$. The overall contextual similarity between $\mathcal{O}$ and $\mathcal{U}$ was determined by averaging the cxt-sim scores of paired prompt groups across all API mappings. The final similarity is 0.266. This low similarity highlights a significant difference in the contextual characteristics between the prompts from $\mathcal{O}$ and $\mathcal{U}$, influencing the model’s reference decoding direction (cf. Section IV-A2).
Summary 6: The contextual characteristics of the input prompts, such as the defined variables and function calls, contribute a significant influence to the deprecated API usage.
IV-C RQ3: Library Perspective for Status Quo and Root Causes
We conducted a detailed analysis to examine the LLMs’ completions across different libraries.
IV-C1 Status Quo Analysis
We observe how the AUP and DUR metrics vary with different libraries.
API Usage Plausibility. The results of API usage plausibility are illustrated in the scatter plots depicted in Figure 4, where each data point signifies the AUP of an LLM for a specific library. Across the 8 libraries, most LLMs exhibited relatively low API usage plausibility for both $\mathcal{O}$ and $\mathcal{U}$, with AUP below 30%. Notable exceptions were Pandas and TensorFlow, where CodeLlama-7b, CodeGen-6b, and CodeGen-2b (represented by symbols “ ”, “ ”, and “ ”, respectively) achieved better AUP (around or greater than 40%). This aligns with the results presented in Model Perspective, where CodeLlama-7b, CodeGen-6b, and CodeGen-2b achieved the best results for the overall dataset. On the other hand, among the libraries, completion prompts from SciPy and seaborn posed the most difficulty for LLMs in predicting plausible API usages, with AUP consistently below 15%.
Figure 4: AUP by Different LLMs across Eight Libraries
Summary 7: Most LLMs exhibited relatively low API usage plausibility across the 8 libraries, with AUP below 30%. CodeLlama-7b, CodeGen-6b, and CodeGen-2b achieved better AUP for Pandas (about 45%-65%) and TensorFlow (around 40%).
Deprecated Usage Rate. The results are presented in the scatter plots shown in Figure 5, where each data point represents the DUR of a particular LLM for a specific library. The results reveal significant differences in the usage of deprecated APIs across the libraries, with DUR ranging from approximately 0% to 100%. Specifically, LLMs generally showed low DUR (around or below 20%) for Numpy, scikit-learn, and TensorFlow. In contrast, LLMs exhibited consistently high DUR for SciPy (approximately 30%-50%) and PyTorch (approximately 50%-70%), and unstable DUR for Pandas (approximately 30%-80%), seaborn (approximately 20%-60%), and Transformers (approximately 40%-100%).
Figure 5: DUR by Different LLMs across Eight Libraries
Summary 8: LLMs showed significant differences in the usage of deprecated APIs across various libraries, with DUR ranging from approximately 0% to 100%. LLMs exhibited consistently high DUR for SciPy and PyTorch, and unstable DUR for Pandas and Transformers.
IV-C2 Potential Cause Discussion
After analyzing the completion prompts and LLMs’ predictions, we found that the AUP differences between these libraries were primarily due to the characteristics of the API usage context. More specifically, the usage contexts of certain APIs followed common patterns and were surrounded by related APIs, allowing advanced LLMs like CodeLlama-7b to infer the desired API functionality based on the completion prompt. For example, many utility APIs in TensorFlow, such as the deprecated tensorflow.compat.v1.initialize_all_variables and its replacement tensorflow.compat.v1.global_variables_initializer, were often used in recognizable patterns that were easier for LLMs to predict. Conversely, some APIs were used more flexibly in diverse contexts and lacked obvious combinations with other APIs, leading to low AUP for LLMs in predicting such APIs. For instance, many APIs in SciPy, like the deprecated scipy.misc.comb and its replacement scipy.special.comb, can be used in diverse contexts to produce combinations for a data sequence. Since the data sequence can originate from numerous sources and be structured in various ways (e.g., Numpy array and Pandas Series), it is challenging for LLMs to predict these APIs based on the completion prompt without additional hints.
The characteristics of API deprecations during library evolution also significantly impacted the use of deprecated APIs in LLM-based code completions. Some APIs were deprecated due to simple package refactoring, leading to similar usage patterns for the deprecated APIs and their replacements. For example, in the version 1.9.0 release of PyTorch, many APIs for tensor linear algebra were moved from the torch package to the torch.linalg package without changes to the API parameters or usage patterns (e.g., torch.lstsq $\to$ torch.linalg.lstsq). In such cases, LLMs found it more difficult to distinguish between deprecated APIs and their replacements, resulting in more frequent use of deprecated APIs in the predicted completions. To further validate this attribution, we gathered prompts associated with each library from $\mathcal{O}$ and $\mathcal{U}$, denoted as $\mathcal{O}^{l}$ and $\mathcal{U}^{l}$. We then performed feature extraction and calculated the contextual similarity between $\mathcal{O}^{l}$ and $\mathcal{U}^{l}$ following the same approach as in Section IV-B2. A higher contextual similarity indicates that $\mathcal{O}^{l}$ and $\mathcal{U}^{l}$, before and after API deprecation, share more common contextual characteristics and usage patterns. Higher $\mathcal{O}^{l}$-$\mathcal{U}^{l}$ contextual similarity suggests that it is more challenging for LLMs to discern whether to use deprecated or replacing APIs based on the given prompts. Table IV shows the relationship between $\mathcal{O}^{l}$-$\mathcal{U}^{l}$ contextual similarity and DUR ranges across the eight libraries, indicating that the DUR range generally increases with higher $\mathcal{O}^{l}$-$\mathcal{U}^{l}$ contextual similarity.
TABLE IV: Relationship between $\mathcal{O}^{l}$-$\mathcal{U}^{l}$ Contextual Similarity and DUR Range across Eight Libraries
  Library Similarity DUR Range
  Numpy 0.056 $<$ 20%
Pandas 0.193 30%-80%
scikit-learn 0.060 $<$ 20%
SciPy 0.094 30%-50%
seaborn 0.054 20%-60%
TensorFlow 0.078 $<$ 25%
PyTorch 0.126 50%-70%
Transformers 0.672 40%-100%
Summary 9: The characteristics of API deprecations during library evolution significantly impacted the use of deprecated APIs in LLM-based code completions. Minor changes between deprecated APIs and their replacements, such as simple package refactoring, often led to more pronounced issues with deprecated API usages.
V Lightweight Mitigations
Based on the findings regarding the causes of deprecated API usage, we proposed two lightweight mitigation approaches, which can serve as baselines for future investigation.
V-A Motivation
As analyzed in the RQs, the causes of deprecated API usage in LLM-based completions can be attributed to Model Training, Model Inference, Prompt, and Library aspects. In this section, we further explore the feasibility of mitigating deprecated API usage issues from these four perspectives.
During model training, a direct mitigation is to clean up the code containing deprecated APIs from the training corpora. However, this is impractical due to the constant library evolution. New API deprecations would cause repeatedly training corpora rebuilding and model retraining. From the library perspective, we cannot control the evolution of the library and API deprecation. Therefore, mitigation should focus on decoding strategies in model inference and prompt engineering.
V-B Two Lightweight Approaches
Our basic idea is illustrated in Algorithm 1. Given an LLM, its generation process can generally be formulated as follows:
$\mathbb{T}^{O}=\textbf{{LLM}}(\mathbb{T}^{I}),$
where $\mathbb{T}^{I}$ and $\mathbb{T}^{O}$ are the input token sequence and output token sequence, respectively. In the context of LLM-based code completion, $\mathbb{T}^{I}$ corresponds to the input prompt $pmpt$, and $\mathbb{T}^{O}$ corresponds to the predicted completion $comp$ (line 1). When $comp$ contains a deprecated API $dep$ (line 3), we need to perform fixing approaches, i.e., the Fix procedure, to replace $dep$ with the corresponding replacement $rep$ (line 4). Note that during the Contains procedure in line 3, alias resolution is conducted similarly to the process in Section III-B1.
Input: Prompt $pmpt$, API Mappings $\mathcal{M}$
Output: Completion $comp$
1 $comp\leftarrow\textbf{{LLM}}(pmpt)$
2 for $(dep\to rep)\in\mathcal{M}$ do
3       if $\emph{{{Contains}}}(comp,dep)$ then
4             $comp\leftarrow\textbf{{Fix}}(pmpt,comp,dep,rep)$
5             break
6
Algorithm 1 Deprecation-Aware Code Completion
The Fix procedure includes reconstructing input, through either (i) replacing the deprecated API tokens or (ii) inserting additional replacing prompts, and then regenerating output:
  • Approach 1 - ReplaceAPI: Replacing Deprecated API Tokens then Regenerating. As shown in Algorithm 2, ReplaceDep removes the tokens corresponding to $dep$ and any subsequent tokens from $comp$, and then appends the tokens of $rep$ (line 2). This results in a prefix $prefix$ that includes $rep$. The $prefix$ is concatenated with the $pmpt$, and the LLM generates the suffix (line 3), i.e., arguments for $rep$ and the remaining tokens to complete the code line. This concatenation forms the fixed completion $comp^{*}$.
  • Approach 2 - InsertPrompt: Inserting Additional Replacing Prompt then Regenerating. As shown in Algorithm 3, CreateDepPmpt constructs an additional replacing prompt $pmpt^{\prime}$ (line 2), formatted as inline comments to guide the LLM to use $rep$ instead of $dep$. By expressing the replacing instruction into inline comments, the re-written prompt (i.e., $pmpt\oplus pmpt^{\prime}$) can be naturally processed by the LLM to continue writing the code. The replacing prompt is structured as follows, where “{comp}”, “{dep}”, and “{rep}” are placeholders for the original completion, deprecated API, and replacing API, respectively, and “...” represents the indentation to ensure syntax correctness.
    ...# {comp}
    ...# {dep} is deprecated, use {rep} instead and revise the return value and arguments.
    The created $pmpt^{\prime}$ is then concatenated with $pmpt$ and fed into the LLM to generate a new completion $comp^{*}$ (line 3).
1 Procedure Fix($pmpt,comp,dep,rep$):
2       $prefix\leftarrow\textbf{{ReplaceDep}}(comp,dep,rep)$
3       $suffix\leftarrow\textbf{{LLM}}(pmpt\oplus prefix)$
4       $comp^{*}\leftarrow prefix\oplus suffix$
5       return $comp^{*}$
Algorithm 2 Approach 1 - ReplaceAPI
1 Procedure Fix($pmpt,comp,dep,rep$):
2       $pmpt^{\prime}\leftarrow\textbf{{CreateRepPmpt}}(comp,dep,rep)$
3       $comp^{*}\leftarrow\textbf{{LLM}}(pmpt\oplus pmpt^{\prime})$
4       return $comp^{*}$
Algorithm 3 Approach 2 - InsertPrompt
V-C Evaluation
We conducted experiments to address the following research question:
  • RQ4: How effectively can the proposed approaches fix deprecated API usage in completions?
V-C1 Evaluation Procedure
For each LLM, we selected up-to-dated samples from $\mathcal{U}$ where the LLM predicted DepC completions using deprecated APIs, i.e.,
$\mathcal{T}=\{(pmpt,dep\to rep)\in\mathcal{U}:\text{anno}(\textbf{{LLM}}(pmpt))=\emph{DepC}\},$
to constitute the evaluation data. These samples were chosen because they have corresponding ground-truth completions (e.g., the line following the prompt in Figure 2), which are essential for assessing the effectiveness of the proposed fixing approaches. For each sample, we employed the deprecation-aware code completion illustrated in Algorithm 1 with the two fixing approaches ReplaceAPI and InsertPrompt to prompt the LLM to generate a completion. We used the following three metrics to assess their effectiveness:
  • Fixed Rate (FR): This metric indicates the proportion of RepC completions predicted by the fixing approaches.
  • Edit Similarity (ES) [74]: This metric measures the similarity between the predicted completions and the ground-truth completions by analyzing the edit operations required to transform one into the other.
  • Exact Match (EM): This metric calculates the rate of predicted completions that exactly match the ground-truth completions after normalizing the return values of function calls (i.e., replacing each element in return value with “_”).
V-C2 Results and Analysis
The evaluation results of the proposed fixing approaches are presented in Table V.
TABLE V: Evaluation Results of Proposed Approaches
  Model ReplaceAPI InsertPrompt
FR (%) ES (%) EM (%) FR (%) ES (%) EM (%)
  CodeGen-350m 92.1 82.3 30.8 25.7 58.7 8.9
CodeGen-2b 88.2 84.6 38.8 66.1 66.0 23.3
CodeGen-6b 85.2 85.3 43.5 77.4 72.9 35.3
DeepSeek-1.3b 99.6 80.9 31.7 93.5 77.7 24.4
StarCoder-3b 90.9 85.0 42.2 85.3 72.2 29.1
CodeLlama-7b 99.5 85.7 48.1 95.5 82.0 43.3
GPT-3.5 97.2 76.2 20.5
ReplaceAPI. Using the ReplaceAPI fixing approach (Algorithm 2), all the LLMs achieve high fixed rates (FR), with values exceeding 85%. Failures in fixing are mainly due to syntax errors or incorrect function calls caused by erroneous tokens following the replaced APIs. For example, consider a DepC completion: “meta_graph_def = tf.saved_model.loader.load(…)” predicted by the original completion procedure (Algorithm 1, line 1). ReplaceAPI replaces the deprecated API tf.saved_model.loader.load with its replacement tf.saved_model.load, producing a $prefix$ of “meta_graph_def=tf.saved_model.load” (line 2 of Algorithm 2). However, CodeGen-2b then predicts a $suffix$ of “_meta_graph_def(…)” (line 3 of Algorithm 2), resulting in an erroneous function call: tf.saved_model.load_meta_graph_def(). This issue arises because the replacement operation in ReplaceAPI can disrupt the naturalness of the code context [75] and the LLMs’ decoding process. An additional interesting finding is that for the three versions of CodeGen, the FR decreases as the model size increases, suggesting that larger models might be more sensitive to interventions during decoding.
The completions generated by LLMs using the ReplaceAPI approach exhibit high ES of over 80%, and EM rates between 30% and 50%, with these rates increasing alongside model size. The inaccuracies in the completions often involve incorrect return values and arguments for the replacing APIs. Incorrect return values arise because the replacing API might include different elements compared to the deprecated API, and the ReplaceAPI approach cannot resolve such inconsistencies in the $prefix$ (line 2 of Algorithm 2). The incorrect arguments are primarily due to the LLMs’ limitations in correctly utilizing replacing APIs, especially those with complex argument lists.
InsertPrompt. When applying the InsertPrompt fixing approach (Algorithm 3), the LLMs exhibited significantly varied fixed rates, ranging from 25.7% to 97.2%. This variation suggests that larger models generally possess a stronger capacity to interpret and utilize inserted prompts formatted as inline comments. An exception to this trend is DeepSeek-1.3b, which achieved a notably high FR of 93.5%. This success can be attributed to using the deepseek-coder-1.3b-instruct version, which has robust zero-shot instruction-following capabilities. Moreover, the FR differences among various LLMs highlight their sensitivity to prompt construction [76, 77], which indicates that different LLMs may require specialized additional prompts in InsertPrompt for optimal performance.
Considering edit similarity and exact match, the completions generated by GPT-3.5 and DeepSeek-1.3b using the InsertPrompt approach do not perform as well as their fixed rates suggest. Despite being fine-tuned with an instruct-tuning corpus, they are not specifically fine-tuned on Python code. As a result, they can follow the instructions in the additional prompt to use replacing APIs but struggle to use those APIs correctly.
Comparison. The comparison between ReplaceAPI and InsertPrompt suggests that direct interventions in the decoding process are more effective than zero-shot prompt engineering. However, the results also reveal the potential of InsertPrompt. First, ReplaceAPI cannot be applied to black-box LLMs like GPT-3.5, as their decoding processes cannot be controlled by users. Second, the additional prompt employed by InsertPrompt was not carefully tuned for each LLM and was performed in a zero-shot manner. In the future, fine-tuning the LLMs with instructions specifically designed for fixing deprecated usage could enhance effectiveness.
Summary 10: ReplaceAPI effectively addresses deprecated API usage for all open-source LLMs, achieving fix rates exceeding 85%. The fixed completions demonstrate acceptable accuracy. While InsertPrompt does not achieve sufficient effectiveness and accuracy in fixing completions containing deprecated API usage, it provides future potential of exploration.
VI Discussion
VI-A Implications
Validating LLM-Generated Code Completions and Issuing Alerts for Deprecated API Usages. Our evaluation study reveals that LLMs frequently use deprecated APIs during code completion. Such deprecated API usages can be easily overlooked [25], potentially introducing bugs or security vulnerabilities into software projects. Therefore, implementing a validation mechanism for deprecated API usage in LLM-generated code completions is crucial to ensure the reliability of the code. Such a validation mechanism can be further integrated into the post-processing of LLM-based code completion, such as issuing alerts to developers.
Fixing and Updating Outdated API Knowledge in LLMs by Model-Level Repair. The current fixing approaches mitigate the issues by intervening in decoding process and rewriting prompts, without addressing the outdated knowledge about deprecated API usages stored in the LLMs. Given the constant evolution of libraries, lightweight model repair techniques are potential solutions for fixing outdated knowledge. Model Editing is one such direction, which can be categorized into the following main categories: Memory-based approaches [78, 79, 80], Locating-then-editing approaches [81, 82, 83], and Meta-learning approaches [84, 85]. Compared to the proposed fixing approaches, model editing can directly update the outdated knowledge about deprecated API usages, even incorporating the information of entirely new replacing APIs (i.e., the replacing APIs introduced after model training and unseen in the training corpora) into the LLMs.
Leveraging Retrieval-Augmented Generation for Up-to-dated Code Completion. As discussed in our study findings, a key cause of the deprecated API usage in LLM-based completions is the lack of posterior knowledge about API deprecations. Retrieval-Augmented Generation (RAG) is a suitable technique that can perfectly align with the need for posterior knowledge [86]. We can explore the possibility of adopting RAG to mitigate deprecated API usage in LLM-based code completion by retrieving related knowledge pieces, such as documentation and usage examples.
Designing Agent & Multi-Agent Systems for Incorporating Library Evolution into LLM-Driven Software Development. In modern software development driven by LLMs, the issues brought by library evolution are encountered not only in code completion. With advancements in AI agents [87, 88, 89], we can potentially develop autonomous agents or multi-agent systems capable of automatically discovering deprecated API usages, identifying correct replacements, upgrading dependent libraries, and fixing the code [90]. To ensure comprehensive recognition of deprecated API usage and up-to-date fixes, we should design an effective multi-agent collaboration pipeline. One agent should scan the generated code to identify all pieces related to API usage. Another agent should continuously fetch information online, checking the latest official API documentation to aid in discovery and correction. Finally, a dedicated agent should be responsible for implementing the necessary library upgrades and code fixes.
VI-B Threats to Validity
Internal Threats. The primary threat to our study is the soundness of the static analysis used for function location and result annotation. Given that Python is a dynamic programming language, the lightweight object type resolution and alis resolution we employed may have missed some function calls of deprecated and replacing APIs during the matching process. Nevertheless, they did not contribute bias to deprecated APIs or replacing APIs in our extracted function calls. In the future, we aim to address this issue by implementing advanced type inference techniques. Additionally, our study currently focuses on function-level API deprecation, overlooking parameter-level deprecations. Future research would investigate a broader type of deprecated APIs to provide more comprehensive analysis.
External Threats. A primary threat lies in the choice of Python libraries and LLMs. To mitigate this threat, we reused libraries examined in previous studies and introduced three popular deep learning libraries to ensure diversity and timeliness. For the LLMs, we selected models covering various architectures, model sizes, training corpora, and training strategies to ensure the generalizability of our findings. Another external threat is that the study was conducted solely on the Python language, which may limit the applicability of our findings to other languages such as Java and C#. In the future, we plan to explore the impact of library evolution on LLM-based code completion across a broader range of programming languages.
VII Conclusion
In this work, we conducted an evaluation study to investigate the statuses and causes of deprecated API usages in LLM-based code completion. The study results all evaluated LLMs encounter challenges in predicting plausible API usages and face issues with deprecated API usages, influenced by the distinct code context characteristics of the prompts and the characteristics of API deprecations during the evolution of these libraries. We propose two lightweight fixing approaches to mitigate the deprecated API usages and can serve as baselines for future research. We also provide implications for the research directions for the combination of library evolution and LLM-driven code completion and software development. We released the code and data of our study at the website [91].
Acknowledgement
This research project is supported by the National Key R&D Program of China (2023YFB4503805), the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN), DSO National Laboratories under the AI Singapore Programme (AISG2-GC-2023-008), National Natural Science Foundation of China under Grant No. 62402342, Shanghai Sailing Program (No. 24YF2749500), and the Ministry of Education, Singapore, under its Academic Research Fund Tier 1 (RG96/23).